Suppose a small island nation imports sugar for its population at the world price of $1,500 per ton. The domestic market for sugar is shown below.

2500 2000 1500 1000 500 S World price Domestic price with subsidy 0 4812 16 20 Quantity (tons/day)

If the government provides a subsidy of $500 per ton, then producer surplus will be ______ per day.

Select one:

a. $0

b. $1,000

c. $4,000

d. $8,000
Answer #1

Ans - A) $0

There is no producer surplus at all as the equilibrium quantity at $1500 is 8 tons which shows that the total value of sugar sold = 1500*8= $12000

When the price falls to $1000 due to subsidy of $500,then  the total value of sugar sold =1000*12= $12000

As the total revenue remains the same in both the cases, there is no producer surplus.